# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 2. Preliminaries and Problem Statement
It is easy to see that, for any regular language $L$, the problem $\mathtt{RES}_{\mathrm{bag}}(L)$, hence $\mathtt{RES}_{\mathrm{set}}(L)$, is in NP (see Section 2). {toappendix} We explain here why $\mathtt{RES}_{\mathrm{bag}}(L)$ for $L$ a regular language is always in NP. Note that this implies the same for $\mathtt{RES}_{\mathrm{set}}(L)$. Given a bag database $D$, it suffices to guess a subdatabase $D^{\prime}\subseteq D$, and verify in PTIME that $\sum_{f\in D^{\prime}}\mathtt{mult}(f)$ is small enough and that $D^{\prime}$ is a contingency set, i.e., $D\setminus D^{\prime}\not\models Q$. The latter point, i.e., the inexistence of an $L$-walk in $D\setminus D^{\prime}$, can easily be checked in PTIME by a product construction followed by a reachability test (see, e.g., (DBLP:journals/siamcomp/MendelzonW95, Lemma 3.1)). However, the NP upper bound is not always tight: there are languages $L$ (e.g., $L=a|b$) such that $\mathtt{RES}_{\mathrm{bag}}(L)$ is in PTIME. The goal of our paper is to characterize the languages $L$ for which $\mathtt{RES}_{\mathrm{set}}(L)$ and $\mathtt{RES}_{\mathrm{bag}}(L)$ are in PTIME, and give efficient algorithms for them.
@@
From the recent work of Bodirsky et al. (bodirsky2024complexity) using earlier results on Valued Constraint Satisfaction Problems (VCSP), the following dichotomy on $\mathtt{RES}_{\mathrm{bag}}$ is already known:
Theorem 2.2 ((bodirsky2024complexity, Remark 8.12)).
@@
We point out that this result can be extended to so-called two-way RPQs, and to the case where some relations are declared exogenous (i.e., they cannot be part of a contingency set). Hence, as discussed in the introduction, Theorem 2.2 gives a complete dichotomy across all regular languages, but it also has some shortcomings. (1.) First, the VCSP approach gives no explicit characterization of the class $L$ of tractable languages, or efficient bounds on the complexity of testing whether a given $L$ is PTIME or not: the criterion in (bodirsky2024complexity), using (kolmogorov2019testing), can be computationally checked but takes time exponential in a representation of the dual of $L$. (2.) Second, in tractable cases, the VCSP approach does not give good bounds on the degree of the polynomial (unlike, e.g., explicit reductions to MinCut). We remark that difficulties (1.) and (2.) also occur in works studying other problems where VCSP implies a general dichotomy but without explicit characterizations or tight bounds: this is the case for instance in (hirai2021minimum). (3.) Third, the VCSP approach does not shed light on the combined complexity, i.e., it does not identify cases that are tractable when both the query and database are given as input. (4.) Fourth, the VCSP approach inherently only applies to bag semantics and not to set semantics, so there is no known analogous dichotomy for $\mathtt{RES}_{\mathrm{set}}$. Our results will improve on some of these shortcomings, though unlike (bodirsky2024complexity) they do not establish a full dichotomy.
@@ 3. Tractability for Local Languages
We now start the presentation of our results on the resilience problem by giving our main tractability result in this section. The result applies to local languages [^3] We note that local languages were also recently observed to correspond to a tractable class of RPQs for which probabilistic query evaluation is equivalent to source-to-target reliability, see (amarilli2025approximating). , a well-known subclass of the regular languages (mpri; sakarovitch2009elements), whose definition is recalled below. We show that the resilience problem is tractable for local languages, also in combined complexity, using a specific notion of automata and a product construction to reduce resilience to MinCut. We stress that local languages do not cover all cases of regular languages with tractable resilience: we will show other tractable cases in Section 7. However, in Section 5 we will give a partial converse: resilience is hard for the so-called four-legged languages, which have a specific form of counterexample to being local.
Local languages.
Intuitively, a regular language $L$ is local if membership to $L$ is described by which letter pairs can occur consecutively and which letters can start or can end words. We use an equivalent characterization (mpri, Chapter III, Section 5.1) via so-called local DFAs:
@@
A DFA is local if, for every letter $a\in\Sigma$, all $a$-transitions in the automaton have the same target state. A language $L$ is local if there is a local DFA $A$ that recognizes $L$.
@@
$s_{1}$$s_{2}$$s_{3}$$s_{4}$$s_{5}$$a$$b$$c$$d$$\epsilon$
(c) RO-$\epsilon$NFA $A_{3}$ for $ab|ad|cd$
Figure 2. Examples of local automata (Example 3.2) and RO-$\epsilon$NFAs (see Example 3.2 and Example 3.8). Initial states are indicated by incoming arrows, final states are doubly circled.
The main result of this section is that $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME when $L$ is local:
Theorem 3.3 (Resilience for local languages).
If $L$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is in PTIME. Moreover, the problem is also tractable in combined complexity: given an $\epsilon$NFA $A$ that recognizes a local language $L$, and a database $D$, we can compute $\mathtt{RES}_{\mathrm{bag}}(Q_{L},D)$ in $\tilde{O}(|A|\times|D|\times|\Sigma|)$.
Note that, in the combined complexity upper bound, the input $A$ is not necessarily a local DFA: it suffices to be given an arbitrary $\epsilon$NFA $A$ recognizing $L$, together with the promise that $L$ is local.
Before proving Theorem 3.3, we remark that it implies a PTIME bound on $\mathtt{RES}_{\mathrm{bag}}(L)$ whenever $\mathtt{reduce}(L)$ is local (as $\mathtt{RES}_{\mathrm{bag}}(L)$ is equivalent to $\mathtt{RES}_{\mathrm{bag}}(\mathtt{reduce}(L))$). For instance, $\mathtt{RES}_{\mathrm{bag}}(L_{0})$ for $L_{0}=\{a,aa\}$ is in PTIME because $\mathtt{reduce}(L_{0})=\{a\}$ is local, even though $L_{0}$ is not. In fact, to see if Theorem 3.3 applies to a language $L$, we can always focus on $\mathtt{reduce}(L)$, for the following reason:
{toappendix}
3.1. Preparatory Lemmas
Before proving the results in Section 3, we prove a few lemmas that use the notions of RO-$\epsilon$NFA (defined later in Section 3) and of letter-Cartesian languages (defined later in Definition 5.1). Specifically, we prove that the formalisms of local DFAs (Definition 3.1), of RO-$\epsilon$NFA, and of letter-Cartesian languages, are indeed equivalent. In the process, we give a general lemma about converting $\epsilon$-NFAs to a certain RO-$\epsilon$NFAs (Lemma 3.4) which we will need in some subsequent proofs.
3.1.1. Equivalence Between RO-$\epsilon$NFA and Local DFAs
First, we show that local DFAs and RO-$\epsilon$NFAs recognize the same languages. We show both directions in turn.
{claim}
Let $A$ be a local DFA. Then there is a RO-$\epsilon$NFA $A^{\prime}$ such that $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$.
Proof.
Let $A=(S,I,F,\Delta)$ be the local DFA. We build from it an RO-$\epsilon$NFA $A^{\prime}=(S^{\prime},I,F,\Delta^{\prime})$ defined as follows. We let $S^{\prime}\coloneq S\cup\{s^{\prime}_{a}\mid a\in\Sigma\}$ where the states of the form $s_{a}^{\prime}$ are fresh. For every letter $a\in\Sigma$ that has a transition in $A$, letting $s_{a}$ be the common target state of all $a$-transitions in $A$, we create one $a$-transition $(s_{a}^{\prime},a,s_{a})\in\Delta^{\prime}$, and add $\epsilon$-transitions $(s,\epsilon,s_{a}^{\prime})\in\Delta^{\prime}$ from every state $s\in S$ that has an outgoing $a$-transition in $A$. The resulting automaton $A^{\prime}$ is an RO-$\epsilon$NFA by construction, and there is a clear bijection between the runs of $A$ and those of $A^{\prime}$. ∎
{claim}
Let $A$ be a RO-$\epsilon$NFA. Then there is a local DFA $A^{\prime}$ such that $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$.
Proof.
Let $A$ be the RO-$\epsilon$NFA. We remove $\epsilon$-transitions in PTIME using standard techniques: first ensure that every state having an $\epsilon$-path to a final state is final; then for every letter $a\in\Sigma$ occurring in $A$, for every state $s$ having an $\epsilon$-path to the starting state $s_{a}^{\prime}$ of the unique $a$-transition $(s_{a}^{\prime},a,s_{a})$ of $A$, add an $a$-transition from $s$ to $s_{a}$; finally remove all $\epsilon$-transitions. It is clear that the resulting automaton $A^{\prime}$ is a local DFA that recognizes the same language, which concludes the proof. ∎
3.1.2. Equivalence Between Local Languages and Letter-Cartesian Languages
Next, we observe that local languages are always letter-Cartesian. We first show one implication:
{claim}
@@ 3.1.2. Equivalence Between Local Languages and Letter-Cartesian Languages
Let $L$ be a local DFA recognizing $L$. Let $x\in\Sigma$, and let $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ such that $\alpha x\beta\in L$ and $\gamma x\delta\in L$, and let us show that $\alpha x\delta$ is in $L$ as well. Because $L$ is local, the state $q$ that is reached after reading $\alpha x$ in the accepting run of $\alpha x\beta$ is the same state that is reached after reading $\gamma x$ in the accepting run of $\gamma x\delta$. We can thus compose a run from the initial state to $q$ reading $\alpha x$, with the run from $q$ to some final state reading $\delta$, to form an accepting run for $\alpha x\delta$. This concludes the proof. ∎
To show the second implication, let us give a lemma that encapsulates a construction of a RO-$\epsilon$NFA from an $\epsilon$NFA. We will use this construction in subsequent proofs:
Lemma 3.4.
Given an $\epsilon$NFA $A$, we can build in time $O(|A|\times|\Sigma|)$ a RO-$\epsilon$NFA $A^{\prime}$ having the following properties:
  1. (1)
    $\mathcal{L}(A)\subseteq\mathcal{L}(A^{\prime})$;
  2. (2)
    $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$ holds if and only if $\mathcal{L}(A)$ is letter-Cartesian.
To prove this result, we need some standard definitions on automata (sakarovitch2009elements, Section 1.3.1):
@@
Let $A=(S,I,F,\Delta)$ be an $\epsilon$NFA. We say that a state $s\in S$ is accessible if there is a path in $A$ from a state of $I$ to $s$ using transitions of $\Delta$. We say that $s$ is co-accessible if there is a path in $A$ from $s$ to a state of $F$ using transitions of $\Delta$. We say that $s$ is useful if it is both accessible and coaccessible. We say that $A$ is trimmed (or trim (sakarovitch2009elements)) if every state of $A$ is useful.
A standard result on automata (sakarovitch2009elements) is that they can be trimmed in linear time:
{claim}
Given an $\epsilon$NFA $A$, we can compute in time $O(|A|)$ an $\epsilon$NFA $A^{\prime}$ which is trimmed and such that $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$.
Proof.
We first do one graph traversal in $A$ from $I$ to compute the accessible states in $O(|A|)$. Then, we do one graph traversal from $F$ following the transitions of $\Delta$ backwards to compute the co-accessible states in $O(|A|)$. Then, we obtain $A^{\prime}$ by simply removing all states of $A$ that are not useful, and removing all transitions that involve these states. The computation of $A^{\prime}$ from $A$ is in time $O(|A|)$ overall.
It is clear that $A^{\prime}$ is then trimmed, because every remaining state $s$ is a state which was useful in $A$, and the paths witnessing it still witness that $s$ is useful in $A^{\prime}$. Further, $A^{\prime}$ recognizes the same language as $A$: the construction clearly ensures $\mathcal{L}(A^{\prime})\subseteq\mathcal{L}(A)$, and conversely there are no accepting runs in $A$ involving states that are not useful, so every accepting run in $A$ is also an accepting run of $A^{\prime}$. This concludes the proof. ∎
We can now show Lemma 3.4:
Proof of Lemma 3.4.
Let $A=(S,I,F,\Delta)$ be an $\epsilon$NFA. Write $L\coloneq\mathcal{L}(A)$. We assume by Section 3.1.2 that $A$ has been trimmed in linear time. Let us compute from $A$ the following information:
  • The set $\Sigma_{\mathrm{start}}$ of letters that can start words of $L$, i.e., the letters $a\in\Sigma$ such that there exists a word $a\alpha\in L$ for some $\alpha\in\Sigma^{*}$. This can be computed in linear time by performing a graph traversal from the initial states following the $\epsilon$-transitions, and then considering every non-$\epsilon$ transition that leaves from such states and letting $\Sigma_{\mathrm{start}}$ be the set of the letters that label such transitions.
  • The set $\Sigma_{\mathrm{end}}$ of letters that can end words of $L$, i.e., letters $a\in\Sigma$ such that $\alpha a\in L$ for some $\alpha\in\Sigma^{*}$. These can be computed in linear time in a symmetric way by doing a graph traversal from $F$, traversing $\epsilon$-transitions backwards, and taking the letters of non-$\epsilon$ transitions leaving backwards from the states reached.
  • The set $\Pi\subseteq\Sigma^{2}$ of pairs of letters that can occur consecutively in a word of $L$, formally, the pairs $(a,b)$ such that there exist $\alpha,\beta\in\Sigma^{*}$ such that $\alpha ab\beta\in L$. We compute this in time $O(|\Sigma|\times|A|)$ by taking each $a\in\Sigma$ and computing the $b\in\Sigma$ such that $(a,b)\in\Pi$. For this, we take the set of the states that are heads of $a$-transitions, and we do a linear-time graph traversal from the set of such states, following $\epsilon$-transitions, and taking all letters that label transitions that leave from states reached in the traversal.
  • The Boolean information of whether $A$ accepts the empty word, which can be computed in linear time by performing a graph traversal from the set $I$ of initial states of $A$ following the $\epsilon$-transitions and checking whether some final state of $F$ is reached.
Note that this uses the fact that the automaton is trimmed. We now define an RO-$\epsilon$NFA $A^{\prime}$ in the following way:
  1. (1)
    For every $a\in\Sigma$, create two states $s_{a}^{\mathtt{in}}$ and $s_{a}^{\mathtt{out}}$, with $(s_{a}^{\mathtt{in}},a,s_{a}^{\mathtt{out}})\in\Delta$.
  2. (2)
    For every $a\in\Sigma_{\mathrm{start}}$, make $s_{a}^{\mathtt{in}}$ an initial state.
  3. (3)
    For every $(a,b)\in\Pi$ (potentially $a=b$), add the $\epsilon$-transition $(s_{a}^{\mathtt{out}},\epsilon,s_{b}^{\mathtt{in}})$ to $\Delta$.
  4. (4)
    For every $a\in\Sigma_{\mathrm{end}}$, make $s_{a}^{\mathtt{out}}$ a final state.
  5. (5)
    If $\epsilon\in L$, create an isolated state that is both initial and final.
It is clear by definition that $A^{\prime}$ is a RO-$\epsilon$-NFA. Write $L^{\prime}\coloneq\mathcal{L}(A^{\prime})$. What remains to show is that $L\subseteq L^{\prime}$, and that we have $L=L^{\prime}$ iff $L$ is letter-Cartesian.
We start by showing that $L\subseteq L^{\prime}$. Let $\alpha\in L$. If $\alpha=\epsilon$ then $\alpha\in L^{\prime}$ by construction. Otherwise let $\alpha=a_{1}\cdots a_{m}$ with $m\geq 1$. The word $\alpha$ witnesses the following:
  • We have $a_{1}\in\Sigma_{\mathrm{start}}$;
  • We have $a_{m}\in\Sigma_{\mathrm{end}}$;
  • We have $(a_{i},a_{i+1})\in\Pi$ for each $1\leq i<m$.
Thus, by construction:
  • The state $s_{a_{1}}^{\mathtt{in}}$ is initial;
  • The state $s_{a_{m}}^{\mathtt{out}}$ is final;
  • We have $(s_{a_{i}}^{\mathtt{out}},\epsilon,s_{a_{i+1}}^{\mathtt{in}})\in\Delta$ for each $1\leq i<m$.
All of this gives an accepting run of $A^{\prime}$ on $\alpha$, hence $\alpha\in L^{\prime}$ indeed.
Now, let us show the equivalence: $L$ is letter-Cartesian iff $L^{\prime}=L$. One direction is easy: if $L^{\prime}=L$, then $L$ is recognized by $A^{\prime}$, which is a RO-$\epsilon$NFA, and thus recognizes a letter-Cartesian language by Section 3.1.1 and Section 3.1.2.
For the converse direction, let us assume that $L$ is letter-Cartesian, and show that $L^{\prime}=L$. The inclusion $L\subseteq L^{\prime}$ was already shown above, so it suffices to show that $L^{\prime}\subseteq L$. Take $\alpha$ in $L^{\prime}$, let us show that it is in $L$. If $\alpha$ is the empty word then it is clear by construction that $\alpha$ is also in $L$. So assume $\alpha=a_{1}\cdots a_{m}$ with $m\geq 1$. By construction of $A^{\prime}$, the fact that $\alpha$ is accepted witnesses that:
  • The state $s_{a_{1}}^{\mathtt{in}}$ is initial
  • The state $s_{a_{m}}^{\mathtt{out}}$ is final;
  • There is an $\epsilon$-transition between each $s_{a_{i}}^{\mathtt{out}}$ and $s_{a_{i+1}}^{\mathtt{in}}$ for each $1\leq i<m$
Therefore, by construction, we have:
@@
  • $a_{1}\in\Sigma_{\mathrm{start}}$, hence there exists $\alpha^{\prime\prime}\Sigma^{*}$ such that $a_{1}\alpha^{\prime\prime}\in L$;
@@
  • $a_{m}\in\Sigma_{\mathrm{end}}$ hence there exists $\alpha^{\prime\prime\prime}\Sigma^{*}$ such that $\alpha^{\prime\prime\prime}a_{m}\in L$;
@@
To do this, we show by finite induction that $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}\alpha^{\prime\prime}_{i}\in L$ for each $1\leq i<m$. The base case of $\alpha^{\prime}_{1}a_{1}a_{2}\alpha^{\prime\prime}_{1}\in L$ is by the definition of $\alpha_{1}^{\prime}$ and $\alpha_{1}^{\prime\prime}$ above. The induction case for $11<i<m-$ 1 is as follows: from $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}\alpha^{\prime\prime}_{i}\in L$ (by induction hypothesis) and $\alpha_{i+1}^{\prime}a_{i+1}a_{i+2}\alpha^{\prime\prime}_{i+1}\in L$ (by definition of $\alpha_{i+1}^{\prime},\alpha_{i+1}^{\prime\prime}$ above), we deduce that $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}a_{i+2}\alpha^{\prime\prime}_{i+1}\in L$ because $L$ is letter-Cartesian on $a_{i+1}$, which is what we wanted to show to establish the induction case. Hence, the inductive claim holds and for $i=m-1$ we deduce that $\alpha_{1}^{\prime}a_{1}\cdots a_{m}\alpha^{\prime\prime}_{m-1}\in L$, which is claim (*). Now, using that $L$ is letter-Cartesian on $a_{1}$, from (*) and the word $a_{1}\alpha^{\prime\prime}\in L$ identified earlier, we deduce that $\alpha\alpha^{\prime\prime}_{m-1}$ is in $L$. Symmetrically, from this and the word $\alpha^{\prime\prime\prime}a_{m-1}\in L$ identified earlier, using the fact that $L$ is letter-Cartesian on $a_{m-1}$ we deduce that $\alpha\in L$, which concludes. ∎
From the lemma, we can now show the converse direction to Section 3.1.2, giving a self-contained proof of the following known result:
Proposition 3.6 ((sakarovitch2009elements, Chp. II, Ex. 1.8, (d))).
A language is local if, and only if, it is letter-Cartesian.
@@
The “only if” direction is shown in Section 3.1.2, so let us focus on the “if” direction. Let then $L$ be a letter-Cartesian language. Let us define directly from $L$ the sets $\Sigma_{\mathrm{start}}$, $\Sigma_{\mathrm{end}}$, and $\Pi$ used in the proof of Lemma 3.4. (Note that we do not know yet that $L$ is regular, so we cannot take an $\epsilon$NFA recognizing $L$ and invoke Lemma 3.4 directly; we must instead define the sets directly from $L$. Further note that, in the statement of the present proposition, we do not talk about the complexity of computing these sets, and do not specify how $L$ is represented.)
From the sets $\Sigma_{\mathrm{start}}$, $\Sigma_{\mathrm{end}}$, and $\Pi$, let us build the RO-$\epsilon$NFA $A^{\prime}$ as in the proof of Lemma 3.4. One can check that the same reasoning as in the proof of Lemma 3.4 implies that, because $L$ is letter-Cartesian, we have $L=\mathcal{L}(A^{\prime})$. This allows us to conclude by Section 3.1.1 that $L$ is local. ∎
3.2. Proof of Section 3.2: Reduction Preserves Locality
{lemmarep}
[Reduction preserves locality] If $L$ is local then so is $\mathtt{reduce}(L)$.

### v3

@@ 2. Preliminaries and Problem Statement
It is easy to see that, for any regular language $L$, the problem $\mathtt{RES}_{\mathrm{bag}}(L)$, hence $\mathtt{RES}_{\mathrm{set}}(L)$, is in NP. Indeed, given a bag database $D$, it suffices to guess a subdatabase $D^{\prime}\subseteq D$, and verify in PTIME that $\sum_{f\in D^{\prime}}\mathtt{mult}(f)$ is small enough and that $D^{\prime}$ is a contingency set, i.e., $D\setminus D^{\prime}\not\models Q$. The latter point, i.e., the nonexistence of an $L$-walk in $D\setminus D^{\prime}$, can easily be checked in PTIME by a product construction followed by a reachability test (see, e.g., (DBLP:journals/siamcomp/MendelzonW95, , Lemma 3.1)).
However, the NP upper bound is not always tight: there are languages $L$ (e.g., $L=a|b$) such that $\mathtt{RES}_{\mathrm{bag}}(L)$ is in PTIME. The goal of our paper is to characterize the languages $L$ for which $\mathtt{RES}_{\mathrm{set}}(L)$ and $\mathtt{RES}_{\mathrm{bag}}(L)$ are in PTIME, and give efficient algorithms for them.
@@
The recent work of Bodirsky et al. (bodirsky2024complexity, ) uses earlier results on Valued Constraint Satisfaction Problems (VCSP) to obtain the following dichotomy on $\mathtt{RES}_{\mathrm{bag}}$:
Theorem 2.2 ((bodirsky2024complexity, , Remark 8.12)).
@@
We point out that this result can be extended to so-called two-way RPQs, and to the case where some relations are declared exogenous (i.e., they cannot be part of a contingency set). Hence, as discussed in the introduction, Theorem 2.2 gives a complete dichotomy across all regular languages. However, it also has four important shortcomings: 1  The VCSP approach gives no explicit characterization of the class $L$ of tractable languages, or efficient bounds on the complexity of testing whether a given $L$ is PTIME or not: the criterion in (bodirsky2024complexity, ), using (kolmogorov2019testing, ), can be computationally checked but takes time exponential in a representation of the so-called dual of $L$. 2  In tractable cases, the VCSP approach does not give good bounds on the degree of the polynomial (unlike, e.g., explicit reductions to MinCut). We remark that the first two difficulties also occur in works studying other problems where VCSP implies a general dichotomy but without explicit characterizations or tight bounds, for instance in (hirai2021minimum, ). 3  The VCSP approach does not shed light on the combined complexity, i.e., it does not identify cases that are tractable when both the query and the database are given as input. 4  The VCSP approach inherently only applies to bag semantics and not to set semantics, so there is no known analogous dichotomy for $\mathtt{RES}_{\mathrm{set}}$. Our results improve on some of these shortcomings. Specifically, all our intractability results already apply to the $\mathtt{RES}_{\mathrm{set}}$ problem, they apply to combined complexity (when the query is given as an automaton along with the database), and they feature a polynomial complexity with a low degree (and in fact up to polylogarithmic factors the data complexity is always linear except for one result, Proposition 7.6, where it is quadratic). However, unlike (bodirsky2024complexity, ), our results do not establish a full dichotomy.
@@ 3. Tractability for Local Languages
We now start the presentation of our results on the resilience problem by giving our main tractability result in this section. The result applies to local languages [^3] We note that local languages were also recently observed to correspond to a tractable class of RPQs for which probabilistic query evaluation is equivalent to source-to-target reliability, see (amarilli2025approximating, ). , a well-known subclass of the regular languages (mpri, ; sakarovitch2009elements, ), whose definition we recall below. We show that the resilience problem is tractable for local languages, also in combined complexity, using a specific notion of automata and a product construction to reduce resilience to MinCut. We stress that local languages do not cover all cases of regular languages with tractable resilience: we will show other tractable cases in Section 7. However, in Section 5 we will give a partial converse: resilience is hard for the so-called four-legged languages, which have a specific form of counterexample to being local.
The remainder of this section is structured as follows: in Section 3.1 we define what are local languages and provide an equivalent characterization that we will use later. We then show in Section 3.2 that resilience is PTIME for local languages.
3.1. Local Languages
We define here local languages and show some useful properties. Intuitively, a language $L$ is local if membership to $L$ is described by which letter pairs can occur consecutively and which letters can start or can end words. To define them formally, we use the notion of so-called local DFAs (mpri, , Chapter III, Section 5.1):
@@ 3.1. Local Languages
A DFA is local if, for every letter $a\in\Sigma$, all $a$-transitions in the automaton have the same target state. A language $L$ is local if there exists a local DFA $A$ that recognizes $L$.
@@
[…]
Letter-Cartesian languages.
We give a known equivalent language-theoretic characterization of local languages, that we call being letter-Cartesian and that will be useful later:
Definition 3.3 (Letter-Cartesian).
Note that, by exchanging the roles of $\alpha,\beta$ and $\gamma,\delta$, if $\alpha x\beta\in L$ and $\gamma x\delta\in L$ then also $\gamma x\beta\in L$. Hence the name letter-Cartesian: for each letter $x\in\Sigma$, the words of $L$ containing an $x$ can be expressed as a Cartesian product between what can appear left of an $x$ and what can appear right of an $x$. Formally, we have $L=\bigcup_{x\in\Sigma}L_{x}xR_{x}$ for some $L_{x},R_{x}\subseteq\Sigma^{*}$.
Example 3.4.
The local languages of Example 3.2 are letter-Cartesian. Take now $L=aa$. With $x\coloneq a$, $\alpha\coloneq a$, $\beta\coloneq\epsilon$, $\gamma\coloneq\epsilon$, $\delta\coloneq a$, we see that $L$ is not letter-Cartesian (as $aaa\notin L$), and hence, as we will see, it is not local.
It is known that letter-Cartesian languages are precisely the local languages:
Proposition 3.5.
A language is local if, and only if, it is letter-Cartesian.
This result has been claimed in (sakarovitch2009elements, , Chp. II, Ex. 1.8, (d)) without proof (as an exercise). For completeness, and to provide more intuition on local languages to the reader, we provide a proof here. One direction is easy to show: local languages are always letter-Cartesian.
Claim 3.6.
@@
Let $A$ be a local DFA recognizing $L$. Let $x\in\Sigma$, and let $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ such that $\alpha x\beta\in L$ and $\gamma x\delta\in L$, and let us show that $\alpha x\delta$ is in $L$ as well. Because $L$ is local, the state $q$ that is reached in $A$ after reading $\alpha x$ in the accepting run of $\alpha x\beta$ is the same state that is reached after reading $\gamma x$ in the accepting run of $\gamma x\delta$. We can thus compose a run from the initial state to $q$ reading $\alpha x$, with the run from $q$ to some final state reading $\delta$, to form an accepting run for $\alpha x\delta$. This concludes the proof. ∎
We now prove the converse, thus establishing Proposition 3.5:
Claim 3.7.
If $L$ is letter-Cartesian then it is local.
To establish this claim, we use a construction from (berstel1996local, , Proposition 2.1). We define the local overapproximation of a language $L$ as the local DFA constructed as follows:
@@
Let $L$ be a language. We start by defining the following:
@@
  • The set $\Sigma_{\mathrm{start}}$ of letters that can start words of $L$, i.e., the letters $a\in\Sigma$ such that there exists a word $a\alpha\in L$ for some $\alpha\in\Sigma^{*}$.
@@
  • The set $\Sigma_{\mathrm{end}}$ of letters that can end words of $L$, i.e., letters $a\in\Sigma$ such that $\alpha a\in L$ for some $\alpha\in\Sigma^{*}$.
  • The set $\Pi\subseteq\Sigma^{2}$ of pairs of letters that can occur consecutively in a word of $L$, formally, the pairs $(a,b)$ such that there exist $\alpha,\beta\in\Sigma^{*}$ such that $\alpha ab\beta\in L$.
We now define the DFA $A$ in the following way. We have an initial state $q_{0}$, which is final iff $\epsilon\in L$. For every $a\in\Sigma$, we have a state $q_{a}$, which is final iff $a\in\Sigma_{\mathrm{end}}$. Then, for every $a\in\Sigma_{\mathrm{start}}$ we add an $a$-transition from $q_{0}$ to $q_{a}$, and for every $(a,b)\in\Pi$ we add a $b$-transition from $q_{a}$ to $q_{b}$.
The DFA $A$ built from $L$ is local by definition, and the following is immediate:
Claim 3.9.
We have $\mathcal{L}(A)\supseteq L$.
We now claim that, for letter-Cartesian languages, the local overapproximation is in fact equal to the original language, namely:
Claim 3.10.
If $L$ is letter-Cartesian, then $\mathcal{L}(A)=L$.
This immediately implies 3.7: if $L$ is letter-Cartesian, then it is local, because $L$ is equal to $\mathcal{L}(A)$ for $A$ the local overapproximation, and $A$ is a local DFA. So all that remains is to show 3.10.
Proof of 3.10.
We must show that if $L$ is letter-Cartesian then $\mathcal{L}(A)=L$. By 3.9, it suffices to show that $\mathcal{L}(A)\subseteq L$.
Take $\alpha$ in $\mathcal{L}(A)$, and let us show that it is in $L$. If $\alpha$ is the empty word then it is clear by construction that $\alpha$ is also in $L$. So assume $\alpha=a_{1}\cdots a_{m}$ with $m\geq 1$. By construction of $A$, the fact that $\alpha$ is accepted witnesses that:
  • $a_{1}\in\Sigma_{\mathrm{start}}$, hence there exists $\alpha^{\prime\prime}\in\Sigma^{*}$ such that $a_{1}\alpha^{\prime\prime}\in L$;
  • $a_{m}\in\Sigma_{\mathrm{end}}$ hence there exists $\alpha^{\prime\prime\prime}\in\Sigma^{*}$ such that $\alpha^{\prime\prime\prime}a_{m}\in L$;
@@
To do this, we show by finite induction that $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}\alpha^{\prime\prime}_{i}\in L$ for each $1\leq i<m$. The base case of $\alpha^{\prime}_{1}a_{1}a_{2}\alpha^{\prime\prime}_{1}\in L$ is by the definition of $\alpha_{1}^{\prime}$ and $\alpha_{1}^{\prime\prime}$ above. The induction case for 1<i<m11<i<m-1 is as follows: from $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}\alpha^{\prime\prime}_{i}\in L$ (by induction hypothesis) and $\alpha_{i+1}^{\prime}a_{i+1}a_{i+2}\alpha^{\prime\prime}_{i+1}\in L$ (by definition of $\alpha_{i+1}^{\prime},\alpha_{i+1}^{\prime\prime}$ above), we deduce that $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}a_{i+2}\alpha^{\prime\prime}_{i+1}\in L$ because $L$ is letter-Cartesian on $a_{i+1}$, which is what we wanted to show to establish the induction case. Hence, the inductive claim holds and for i=m1i=m-1 we deduce that $\alpha_{1}^{\prime}a_{1}\cdots a_{m}\alpha^{\prime\prime}_{m-1}\in L$, which is claim (*). Now, using that $L$ is letter-Cartesian on $a_{1}$, from (*) and the word $a_{1}\alpha^{\prime\prime}\in L$ identified earlier, we deduce that $\alpha\alpha^{\prime\prime}_{m-1}$ is in $L$. Symmetrically, from this and the word $\alpha^{\prime\prime\prime}a_{m}\in L$ identified earlier, using the fact that $L$ is letter-Cartesian on $a_{m}$ we deduce that $\alpha\in L$, which concludes. ∎
We have concluded the proof of Proposition 3.5.
Testing if a language is local.
We can use the local overapproximation (Definition 3.8) to show that it is PTIME to determine whether a language is local. Indeed, the previous results imply:
Claim 3.11.
Let $L$ a language, and $A$ its local overapproximation. Then we have $\mathcal{L}(A)=L$ if and only if $L$ is local.
@@
If $L$ is local, then $L$ is letter-Cartesian by Proposition 3.5 so by 3.10 we have $\mathcal{L}(A)=L$. Conversely, if $\mathcal{L}(A)=L$ then $L$ is recognized by a local DFA, so it is local. ∎
This result gives us a way to check whether the language recognized by an automaton is local. First, in the case of DFA, we can check this tractably:
Proposition 3.12 (Locality testing for DFAs).
Given a DFA $A$, it is in PTIME to determine if $\mathcal{L}(A)$ is local.